I cam in the noon to check quality of optics runs without sieve. Runs with dp = 0% and dp=-2% looks fine. Example plots are attached. However, since dp = -4%, only HRS momentum was lowered while septum momentum remains at setpoint of dp=-2%. Further check on match between HRS momentum and septum momentum was performed as 3rd plot. There was two mismatch periods: * Jun 27, 6:15AM - 6:30AM: No data was taken; * Jun 27, 10:44AM - noon shudown: dp = -4% and -10% runs have this problem which are not useful for optics calibration. But they are still good for detector check out. In the end, we still have sieved -4% and -10% data to calibrate optics in this region.
Run 1212: dp = 0%
Run 1241: dp = -2%
GREEN: HRS p set point; Blue: septum current